# Duality theory in linear optimization and its extensions — formally verified

### v2

@@ 7.1 Counterexamples for LABEL:extendedFarkas
Recall that LABEL:extendedFarkas has four preconditions on matrix $A$ and vector $b$. The following examples show that omitting any of these preconditions makes the theorem false — there may exist nonnegative vectors $x,y$ such that $A\cdot x\leq b$, $(-A^{T})\cdot y\leq 0$, and $b\cdot y<0$.
$A=\begin{pmatrix}\bot&\top\\ 0&-1\end{pmatrix}\hfil\hskip 30.&b=\begin{pmatrix}0\\ -1\end{pmatrix}\hfil\hskip 60.&x=\begin{pmatrix}1\\ 1\end{pmatrix}\hfil\hskip 30.&y=\begin{pmatrix}0\\ 1\end{pmatrix}\hfil\hskip 60.&\mbox{ $A$ has $\bot$ and $\top$ in the same row}\vspace{12pt}\\ A=\begin{pmatrix}\bot\\ \top\end{pmatrix}\hfil\hskip 30.&b=\begin{pmatrix}-1\\ 0\end{pmatrix}\hfil\hskip 60.&x=\begin{pmatrix}0\end{pmatrix}\hfil\hskip 30.&y=\begin{pmatrix}1\\ 1\end{pmatrix}\hfil\hskip 60.&\mbox{ $A$ has $\bot$ and $\top$ in the same column}\vspace{12pt}\\ A=\begin{pmatrix}\top\\ -1\end{pmatrix}\hfil\hskip 30.&b=\begin{pmatrix}\top\\ -1\end{pmatrix}\hfil\hskip 60.&x=\begin{pmatrix}1\end{pmatrix}\hfil\hskip 30.&y=\begin{pmatrix}0\\ 1\end{pmatrix}\hfil\hskip 60.&\mbox{ $A$ has $\top$ in a row where $b$ has $\top$}\vspace{12pt}\\ A=\begin{pmatrix}\bot\end{pmatrix}\hfil\hskip 30.&b=\begin{pmatrix}\bot\end{pmatrix}\hfil\hskip 60.&x=\begin{pmatrix}1\end{pmatrix}\hfil\hskip 30.&y=\begin{pmatrix}0\end{pmatrix}\hfil\hskip 60.&\mbox{ $A$ has $\bot$ in a row where $b$ has $\bot$}\end{array}\vspace{12pt}\$
@@
$A=\begin{pmatrix}\bot\\ 0\end{pmatrix}\hfil\hskip 30.&b=\begin{pmatrix}0\\ -1\end{pmatrix}\end{array}$
@@
Let us consider the following six LPs over $F_{\infty}$; all of them are written in the format $P=(A,b,c)$.
$P_{1}=\left(\begin{pmatrix}\bot\\ \top\end{pmatrix},\begin{pmatrix}-1\\ 0\end{pmatrix},\begin{pmatrix}0\end{pmatrix}\right)\hfil\hskip 100.&D_{1}=\left(\begin{pmatrix}\top&\bot\end{pmatrix},\begin{pmatrix}0\end{pmatrix},\begin{pmatrix}-1\\ 0\end{pmatrix}\right)\vspace{5pt}\\ P_{2}=\left(\begin{pmatrix}\bot\end{pmatrix},\begin{pmatrix}\bot\end{pmatrix},\begin{pmatrix}0\end{pmatrix}\right)\hfil\hskip 100.&D_{2}=\left(\begin{pmatrix}\top\end{pmatrix},\begin{pmatrix}0\end{pmatrix},\begin{pmatrix}\bot\end{pmatrix}\right)\vspace{5pt}\\ P_{3}=\left(\begin{pmatrix}\top\\ -1\end{pmatrix},\begin{pmatrix}\top\\ -1\end{pmatrix},\begin{pmatrix}0\end{pmatrix}\right)\hfil\hskip 100.&D_{3}=\left(\begin{pmatrix}\bot&1\end{pmatrix},\begin{pmatrix}0\end{pmatrix},\begin{pmatrix}\top\\ -1\end{pmatrix}\right)\end{array}$
It can be checked that $D_{i}$ is the dual of $P_{i}$ for $i=1,2,3$. Furthermore, strong duality fails in all cases, since the optimum of $P_{i}$ is 00 and the optimum of $D_{i}$ is $\bot$ for $i=1,2,3$. Each of $P_{1},D_{1},P_{2},D_{2},P_{3},D_{3}$ violates exactly one of the conditions in the definition of a valid LP.
@@ 8.1 Farkas-like theorems
  • Bottesch, Haslbeck, Thiemann [8] proved a variant of LABEL:equalityFarkas for $\delta$-rationals by analyzing a specific implementation of the Simplex algorithm [18] by Marić, Spasić, Thiemann.
  • Bottesch, Raynaud, Thiemann [9] proved the Fundamental theorem of linear inequalities as well as both LABEL:equalityFarkas and LABEL:inequalityFarkas for all linearly ordered fields, alongside with the Carathéodory’s theorem and the Farkas-Minkowski-Weyl theorem. They also investigated systems of linear mixed-integer inequalities.
  • Thiemann himself [28] then proved the strong LP duality in the asymmetric version (3).
Sakaguchi [26] proved a version of LABEL:equalityFarkas for linearly ordered fields in Rocq, using the Fourier-Motzkin elimination. Allamigeon and Katz [3] made a large contribution to the study of convex polyhedra in Rocq — among other results, they proved a version of LABEL:equalityFarkas for linearly ordered fields as well as the strong LP duality in the asymmetric version (3).
@@ 8.2 Hahn-Banach theorems
Several Hahn-Banach theorems have been formalized in Lean as a part of Mathlib. In particular, it would have been possible to prove LABEL:equalityFarkas for reals using the Hahn-Banach separation theorem for a convex closed set and a point [7]. It would then have been possible to extend the result to other duality theorems for reals. To our knowledge, the theorem [7] cannot be used to prove LABEL:equalityFarkas for an arbitrary linearly ordered field. Therefore, we decided to prove LABEL:equalityFarkas without appealing to the geometry and topology.
Note that certain Hahn-Banach theorems have also been formalized in Mizar [23], Alf [22], Isabelle [6], and Rocq [15].
@@ Acknowledgements
We would like to thank David Bartl and Jasmin Blanchette for frequent consultations. We would also like to express gratitude to Andrew Yang for the proof of Finset.univ_sum_of_zero_when_not and to Henrik Böving for a help with generalization from extended rationals to extended linearly ordered fields. We would also like to acknowledge Antoine Chambert-Loir, Apurva Nakade, Yaël Dillies, Richard Copley, Edward van de Meent, Markus Himmel, Mario Carneiro, and Kevin Buzzard.

### v3

@@ 7.1 Counterexamples for LABEL:extendedFarkas
Recall that LABEL:extendedFarkas has four preconditions on matrix $A$ and vector $b$. The following examples show that omitting any of these preconditions makes the theorem false.
  • If $A$ has $\bot$ and $\top$ in the same row, it may happen that both $x$ and $y$ exist:
    $A=\begin{pmatrix}\bot&\top\\ 0&-1\end{pmatrix}\qquad\qquad b=\begin{pmatrix}0\\ -1\end{pmatrix}\qquad\qquad x=\begin{pmatrix}1\\ 1\end{pmatrix}\qquad\qquad y=\begin{pmatrix}0\\ 1\end{pmatrix}$
  • If $A$ has $\bot$ and $\top$ in the same column, it may happen that both $x$ and $y$ exist:
    $A=\begin{pmatrix}\bot\\ \top\end{pmatrix}\qquad\qquad b=\begin{pmatrix}-1\\ 0\end{pmatrix}\qquad\qquad x=\begin{pmatrix}0\end{pmatrix}\qquad\qquad y=\begin{pmatrix}1\\ 1\end{pmatrix}$
  • If $A$ has $\top$ in a row where $b$ has $\top$, it may happen that both $x$ and $y$ exist:
    $A=\begin{pmatrix}\top\\ -1\end{pmatrix}\qquad\qquad b=\begin{pmatrix}\top\\ -1\end{pmatrix}\qquad\qquad x=\begin{pmatrix}1\end{pmatrix}\qquad\qquad y=\begin{pmatrix}0\\ 1\end{pmatrix}$
  • If $A$ has $\bot$ in a row where $b$ has $\bot$, it may happen that both $x$ and $y$ exist:
    $A=\begin{pmatrix}\bot\end{pmatrix}\qquad\qquad b=\begin{pmatrix}\bot\end{pmatrix}\qquad\qquad x=\begin{pmatrix}1\end{pmatrix}\qquad\qquad y=\begin{pmatrix}0\end{pmatrix}$
@@
[…]
@@
Let us consider the following six LPs over $F_{\infty}$; all of them are written in the format $(A,b,c)$.
[…]
It can be checked that $D_{i}$ is the dual of $P_{i}$ for $i=1,2,3$. Furthermore, strong duality fails in all cases, since the optimum of $P_{i}$ is 0 and the optimum of $D_{i}$ is $\bot$ for $i=1,2,3$. Each of $P_{1},D_{1},P_{2},D_{2},P_{3},D_{3}$ violates exactly one of the conditions in the definition of a valid LP.
@@ 8.1 Farkas-like theorems
  • Bottesch, Haslbeck, Thiemann [Farkas-AFP] proved a variant of LABEL:equalityFarkas for $\delta$-rationals by analyzing a specific implementation of the Simplex algorithm [Simplex-AFP] by Marić, Spasić, Thiemann.
  • Bottesch, Raynaud, Thiemann [Linear-AFP] proved the Fundamental theorem of linear inequalities as well as both LABEL:equalityFarkas and LABEL:inequalityFarkas for all linearly ordered fields, alongside with the Carathéodory’s theorem and the Farkas-Minkowski-Weyl theorem. They also investigated systems of linear mixed-integer inequalities.
  • Thiemann himself [Duality-AFP] then proved the strong LP duality in the asymmetric version (3).
Sakaguchi [^5] https://github.com/pi8027/vass proved a version of LABEL:equalityFarkas for linearly ordered fields in Rocq, using the Fourier-Motzkin elimination. Allamigeon and Katz [Simplex-Coq] made a large contribution to the study of convex polyhedra in Rocq — among other results, they proved a version of LABEL:equalityFarkas for linearly ordered fields as well as the strong LP duality in the asymmetric version (3).
@@ 8.2 Hahn-Banach theorems
Several Hahn-Banach theorems have been formalized in Lean as a part of Mathlib. In particular, it would have been possible to prove LABEL:equalityFarkas for reals using the Hahn-Banach separation theorem [^6] https://github.com/leanprover-community/mathlib4/blob/master/Mathlib/Analysis/LocallyConvex/Separation.lean for a convex closed set and a point. It would then have been possible to extend the result to other duality theorems for reals. To our knowledge, this theorem however cannot be used to prove LABEL:equalityFarkas for an arbitrary linearly ordered field. Therefore, we decided to prove LABEL:equalityFarkas without appealing to the geometry and topology.
Note that certain Hahn-Banach theorems have also been formalized in Mizar [HahnBanach-Mizar], Alf [HahnBanach-Alf], Isabelle [HahnBanach-Isabelle], and Rocq [HahnBanach-Coq].
@@ Acknowledgements
We would like to thank David Bartl and Jasmin Blanchette for frequent consultations. We would also like to express gratitude to Henrik Böving for a help with generalization from extended rationals to extended linearly ordered fields and to Andrew Yang for the proof of Finset.univ_sum_of_zero_when_not. We would also like to acknowledge Antoine Chambert-Loir, Apurva Nakade, Yaël Dillies, Richard Copley, Edward van de Meent, Markus Himmel, Mario Carneiro, and Kevin Buzzard.